OpenStudy (aurorab):

Choose the correct description of the graph of the compound inequality x - 2 > -4 and 3x <_15 a.A number line with an open circle on -2, shading to the left, and a closed circle on 5, shading to the right. b.A number line with a closed circle on -2, shading to the left, and an open circle on 5, shading to the right c.A number line with an open circle on -2, a closed circle on 5, and shading in between d. A number line with a closed circle on -2, an open circle on 5, and shading in between.

OpenStudy (aurorab):

@vjordillo help plzz

OpenStudy (anonymous):

x - 2 > -4 solve by adding 2 to both sides 3x <_15 solve by dividing both sides by 3 This should make answering your questions more clear

OpenStudy (aurorab):

-4 +2 is -6 right?

OpenStudy (anonymous):

a +2 moves it to the right on a number line a -2 would move it to the left

OpenStudy (anonymous):

+5-5 =0 +5+5=10 -5+5=0 -5-5=-10